Estimate the average force exerted by a shot-putter on a 7kg shot if the shot is moved through a distance of 2.6m and is released with a speed of 10m/s
Answer #1

initial velocity vi=10m/sec

final velocity vf=0m/sec

s=2.6m

thus by equation of motion

vf2-vi2=2as

|a|=vi2/2s= 19.23m/sec2

now by newton's second law

F=ma= 7kgx19.23m/sec2=134.61N
Answer #2

time taken to travel = 2.6/10 = .26sec

force = Ma = 7*10/.26 = 269.23N
Answer #3

v^2-u^2=2*a8s

100=2*a*2.6

a=1.923m/s^2

f=m*a

=7*1.923=13.461 newtons
Answer #4

w = change in K.E ...

F*s = (1/2)*m*(v^2)...
F = (1/2)*m*u^2/s....
F = 134.6 N